What is the force acting on a 5m2 wall by normal atmospheric pressure at sea level (101kPa)?


1
Expert's answer
2022-02-18T08:22:05-0500

Given:

"A=5\:\rm m^2"

"P=101000\:\rm Pa"


The force

"F=pA=101000*5=5.05*10^5\:\rm N"